# Efficient Large-Scale Quantum Optimization via Counterdiabatic Ansatz

### v1

@@ Abstract
Quantum Approximate Optimization Algorithm (QAOA) is one of the fundamental variational quantum algorithms, while a version of QAOA that includes counterdiabatic driving, termed Digitized Counterdiabatic QAOA (DC-QAOA), is generally considered to outperform QAOA for all system sizes when the circuit depth for the two algorithms are held equal. Nevertheless, DC-QAOA introduces more CNOT gates per layer, so the overall circuit complexity is a tradeoff between the number of CNOT gates per layer and the circuit depth, and must therefore be carefully assessed. In this paper, we conduct a comprehensive comparison of DC-QAOA and QAOA on MaxCut problem with the total number of CNOT gates held equal, and we focus on one implementation of counterdiabatic terms using nested commutators in DC-QAOA, termed as DC-QAOA(NC). We have found that DC-QAOA(NC) reduces the overall circuit complexity as compared to QAOA only for sufficiently large problems, and for MaxCut problem the number of qubits must exceed 16 for DC-QAOA(NC) to outperform QAOA. We have further shown that this advantage can be understood from the effective dimensions introduced by the counterdiabatic driving terms. Moreover, based on our finding that the optimal parameters generated by DC-QAOA(NC) strongly concentrate in the parameter space, we haved devised an instance-sequential training method for DC-QAOA(NC) circuits, which, compared to traditional methods, offers performance improvement while using even fewer quantum resources. Our findings provide a more comprehensive understanding of the advantages of DC-QAOA circuits and an efficient training method based on their generalizability.
@@ I INTRODUCTION
In this paper, we compare DC-QAOA and QAOA on the MaxCut problem Aoshima and Iri (1977), where the number of layers is increased until both algorithms reach the predefined fidelity to the ground state of optimization problems. We use the number of CNOT gates as an indicator of efficiency since it characterizes the difficulty of implementing such algorithms on quantum computers. Using MaxCut on randomly generated unweighted graphs as benchmarks, we found that for graphs with over 16 vertices, DC-QAOA can achieve better performance using fewer two-qubit gates, while QAOA is better for smaller graphs. By closely examining the effective dimension of the two circuits, we find that adding cd-terms in each layer provides more effective dimensions than simply adding more layers to QAOA, and the cd-terms can suppress non-diagonal transitions, thereby improving DC-QAOA’s performance. Furthermore, the optimal parameters obtained by training DC-QAOA display a strong concentration in the parameter space, demonstrating potential transferability and generalizability from optimization problems on a smaller scale to those on a larger scale. Based on our findings, we devised an Instance-Sequential Training (IST) method for DC-QAOA, which enables comparable or superior outcomes with substantially reduced quantum resources. As a result, DC-QAOA emerges as a more viable option for extensive quantum optimization tasks, in contrast to QAOA, which may be more efficient for smaller-scale problems.
The remainder of the paper is organized as follows. In Section II, we introduce the formalism including definitions of QAOA and DC-QAOA. Section III provides the results. We show a comprehensive comparision of DC-QAOA and QAOA in Section III.1. In Section III.2, we discuss how DC-QAOA benefits from cd-terms, and in Section III.3, we demonstrate the parameter concentration phenomenon in DC-QAOA and propose a quantum resource-efficient training method. We conclude in Section IV.
@@ II Formalism
where $\lambda(t)\in[0,1]$ is the annealing schedule in $t\in[0,1]$. As $\dot{\lambda}\rightarrow 0$, the ground state of $H_{\mathrm{i}}$ can evolve adiabatically into the ground state of $H_{\mathrm{f}}$. Here, $H_{\mathrm{i}}=\sum_{k}h_{k}\sigma_{k}^{x}$, allowing the ground state of $H_{\mathrm{i}}$ to be easily prepared as an equally weighted superposition state in the computational bases, and $H_{\mathrm{f}}$ can be any Hamiltonian of interest. Quantum annealing can be digitized with trotterized time evolution:
@@
However, the requirement for a large number of trotter layers $p$ in QAOA to approximate adiabatic evolution poses challenges due to the increased circuit complexity. To address this limitation, DC-QAOA has been developed Chandarana et al. (2022). DC-QAOA builds upon analog cd-driving which can shorten the evolution time of adiabatic transitions by adding an extra cd-term, presenting an opportunity to effectively reduce number of layers $p$ in QAOA. The nature of the cd-terms depends on the choice of appropriate adiabatic gauge potential Sels and Polkovnikov (2017). In the realm of analog cd-driving, the commonly utilized gauge potential is first-order nested commutators
@@
A schematic circuit diagram of a typical $p$-layer DC-QAOA is shown in Fig. 1. The detailed implementation of three possible cd-terms in DC-QAOA using elementary gates, including first-order nested commutator (hereafter referred to as NC) Chandarana et al. (2023), single qubit Y-rotation (referred to as Y) Čepaitė et al. (2023), and two-qubit YY-Interaction (referred to as YY) Chai et al. (2022), can be seen in Fig. 2. In this paper, we denote the three versions of the DC-QAOA algorithm which uses NC, YY and Y as cd-terms as DC-QAOA(NC), DC-QAOA(YY), and DC-QAOA(Y), respectively.
@@ III.1 Circuit Length Scaling Behavior
For a MaxCut problem on graph $\mathcal{G}$ having $M$ edges, the implementation of $U_{\mathrm{f}}$ requires $2M$ CNOT gates, and the implementation of $U_{\mathrm{cd}}$ requires $4M$, $2M$ and 00 CNOT gates for NC (Fig. 2(a)), YY (Fig. 2(b)) and Y (Fig. 2(c)), respectively. Note that QAOA has $2M$ CNOT gates, therefore DC-QAOA(NC), DC-QAOA(YY), and DC-QAOA(Y) has $6M$, $4M$, and $2M$ CNOT gates in total respectively. It is worth noting that DC-QAOA(NC) utilizes three times CNOT gates as in QAOA, the most among the three versions of the algorithm.
It has been found that the depth of QAOA required to reach the ground state of Hamiltonian $H_{\mathrm{f}}$ with a predefined fixed probability scales roughly linearly with problem size Akshay et al. (2022). Such depth is called the critical depth, denoted as $p^{*}$, which provides a tool to quantify the advantage afforded by adding cd-terms in DC-QAOA. It is important to see how fast $p^{*}$ scales with problem size. For example, if $p^{*}$ of DC-QAOA(NC) scales three times slower than $p^{*}$ of QAOA, it means that as the problem size grows, DC-QAOA(NC) can eventually solve an optimization problem utilizing fewer CNOT gates than QAOA. Therefore, it is fair to compare DC-QAOA and QAOA by comparing the scaling rate of $p^{*}$ versus the problem size. To find the critical depth $p^{*}$ for DC-QAOA and QAOA, we first define the approximate ratio $R$:
$R=|\braket{\psi_{\mathrm{o}}}{\psi_{\mathrm{g}}}|^{2},$ (9)
which is the fidelity between the output state |$\ket{\psi_{\mathrm{o}}}$ of the quantum algorithm and the ground state |$\ket{\psi_{\mathrm{g}}}$ of $H_{\mathrm{f}}$. Then, we determine the critical depth $p^{*}$ as the minimum number of layers required in DC-QAOA or QAOA for $R$ to exceed 90%.
@@
Fig. 3 shows the approximate ratio $R$, calculated at fixed number of layers $p=20$, for a MaxCut problem with 16 vertices (cf. Fig. 9(h)) versus number of training iterations for four different versions of the algorithms. Different algorithms converge to different values of $R$, with a higher value indicating a better performance. Each panel shows eight color lines, which are 8 runs from random initialization of parameters. The converged value of $R$ for different algorithms are averaged and are shown in Table 1. Overall, the results from DC-QAOA(NC) (Fig. 3(b)) is the best, followed by those from DC-QAOA(YY) (Fig. 3(c)). Results from DC-QAOA(Y) (Fig. 3(d)) and QAOA (Fig. 3(a)) are comparable and are both inferior to the other two methods. These results suggest having more CNOT gates on each layer implies better performance of the algorithm.
@@
Figure 3: The approximate ratio $R$, calculated at fixed number of layers $p=20$, for a MaxCut problem versus number of training iterations for four different versions of the algorithms using same number of layers $p=20$. (a) QAOA, (b) DC-QAOA(NC), (c) DC-QAOA(YY) and (d) DC-QAOA(Y). The target graph has 16 vertices. In each panel, the 8 color lines represents 8 random initialization of parameters.
@@
Table 1: The average approximate ratio $R$ (“Mean”) for four different algorithms, along with their respective standard deviations (“Std.”).
@@
To elucidate how the combined effect of the number of CNOT gates and number of layers on the effectiveness of the algorithm, we focus on comparing DC-QAOA(NC), which has the most number of CNOT gates per layer, to QAOA. In Fig. 4(a) we compare the critical depth $p^{*}$ as functions of the number of qubits $N$, for both QAOA and DC-QAOA(NC). The calculated critical depths for both algorithms scale approximately linearly with system size as $p_{\mathrm{QAOA}}^{*}\simeq 3.48N-20.70$ (for QAOA) and $p_{\mathrm{DC-QAOA(NC)}}^{*}\simeq 0.82N-1.14$ (for DC-QAOA(NC)). Therefore, in the limit of solving large scale MaxCut problem on graph $\mathcal{G}$ with $M$ edges, QAOA needs $3.48N$ layers and in turn $3.48N\times 2M=6.96NM$ CNOT gates and DC-QAOA(NC) needs $0.82N$ layers and in turn $0.82N\times 6M=4.96NM$ CNOT gates. This implies that DC-QAOA(NC) is more advantageous than QAOA in solving large scale optimization problem because the total number of CNOT gates required to achieve the same level of effectiveness is smaller for DC-QAOA(NC), when $N$ is sufficiently large. In other words, the circuit complexity for DC-QAOA(NC) is lower, taking into consideration both the number of effective layers and the number of CNOT gates per layer required.
@@
We set $p_{\mathrm{QAOA}}^{*}=p_{\mathrm{DC-QAOA(NC)}}^{*}$ and obtain $N\simeq 16.94$. This suggests that DC-QAOA(NC) is more advantageous than QAOA when the number of vertices exceeds 16 (i.e. $N\geq 17$). To show this more clearly, we execute DC-QAOA(NC) with $p$ layers, and a QAOA with intentionaly-set $3p$ layers (which is denoted as QAOA-3 in Fig. 4(b),(c) and (d) and should not be confused with regular QAOA with $p$ layers), to solve the MaxCut problem with different number of vertices. Fig. 4(b) shows the results on 12 vertices (cf. Fig. 9(d)), and one can see that QAOA-3 clearly performs better than DC-QAOA(NC). Fig. 4(c) shows the results on 16 vertices (cf. Fig. 9(h)), in which situation the two algorithms are comparable. Given that $N=16$ is still less than 16.94, it is reasonable that QAOA-3 outperforms DC-QAOA(NC) slightly at this stage, although DC-QAOA(NC) rapidly approaches similar performance with increasing $p$. When $p>10$, the difference in $R$ between DC-QAOA(NC) and QAOA becomes negligible. We have also conducted a calculation comparing QAOA-3 and DC-QAOA(NC) for a problem with 20 vertices, with results shown in Fig. 4 (d), although the calculation is more expensive. In this case, DC-QAOA(NC) is clearly superior than QAQA-3. Our results shown in Fig. 4 confirms that DC-QAOA(NC) is only more advantageous than QAOA on graphs with more than 16 vertices. To our knowledge, this fact has not been noted in the literature and is the key result of this paper.
@@
Figure 4: (a) The critical depth $p^{*}$ versus the number of qubits $N$ for DC-QAOA(NC) and QAOA. (b)-(d) The approximate ratio $R$ versus $p$ for DC-QAOA(NC) with $p$ layers and for QAOA with $3p$ layers (QAOA-3) using the same number of CNOT gates to find the maximum cut in a graph with (b) 12, (c) 16 and (d) 20 vertices. The dotted black line represent the predefined value 90%.
III.2 Effective Dimension and Excitation Suppression

### v2

@@ Abstract
Quantum Approximate Optimization Algorithm (QAOA) is one of the fundamental variational quantum algorithms, while a version of QAOA that includes counterdiabatic driving, termed Digitized Counterdiabatic QAOA (DC-QAOA), is generally considered to outperform QAOA for all system sizes when the circuit depth for the two algorithms are held equal. Nevertheless, DC-QAOA introduces more CNOT gates per layer, so the overall circuit complexity is a tradeoff between the number of CNOT gates per layer and the circuit depth and must be carefully assessed. In this paper, we conduct a comprehensive comparison of DC-QAOA and QAOA on MaxCut problem with the total number of CNOT gates held equal, and we focus on one implementation of counterdiabatic terms using nested commutators in DC-QAOA, termed as DC-QAOA(NC). We have found that DC-QAOA(NC) reduces the overall circuit complexity as compared to QAOA only for sufficiently large problems, and for MaxCut problem the number of qubits must exceed 16 for DC-QAOA(NC) to outperform QAOA. Additionally, we benchmark DC-QAOA(NC) against QAOA on the Sherrington-Kirkpatrick model under realistic noise conditions, finding that DC-QAOA(NC) exhibits significantly improved robustness compared to QAOA, maintaining higher fidelity as the problem size scales. Notably, in a direct comparison between one-layer DC-QAOA(NC) and three-layer QAOA where both use the same number of CNOT gates, we identify an exponential performance advantage for DC-QAOA(NC), further signifying its suitability for large-scale quantum optimization tasks. We have further shown that this advantage can be understood from the effective dimensions introduced by the counterdiabatic driving terms. Moreover, based on our finding that the optimal parameters generated by DC-QAOA(NC) strongly concentrate in the parameter space, we have devised an instance-sequential training method for DC-QAOA(NC) circuits, which, compared to traditional methods, offers performance improvement while using even fewer quantum resources. Our findings provide a more comprehensive understanding of the advantages of DC-QAOA circuits and an efficient training method based on their generalizability.
@@ I INTRODUCTION
In this paper, we compare DC-QAOA and QAOA on the MaxCut problem Aoshima and Iri (1977), where the number of layers is increased until both algorithms reach the predefined fidelity to the ground state of optimization problems. We use the number of CNOT gates as an indicator of efficiency since it characterizes the difficulty of implementing such algorithms on quantum computers. Using MaxCut on randomly generated unweighted graphs as benchmarks, we found that for graphs with over 16 vertices, DC-QAOA can achieve better performance using fewer two-qubit gates, while QAOA is better for smaller graphs. By closely examining the effective dimension of the two circuits, we find that adding cd-terms in each layer provides more effective dimensions than simply adding more layers to QAOA, and the cd-terms can reduce non-diagonal transitions, thereby improving DC-QAOA’s performance. Furthermore, the optimal parameters obtained by training DC-QAOA display a strong concentration in the parameter space, demonstrating potential transferability and generalizability from optimization problems on a smaller scale to those on a larger scale. Based on our findings, we devised an Instance-Sequential Training (IST) method for DC-QAOA, which enables comparable or superior outcomes with substantially reduced quantum resources. As a result, DC-QAOA emerges as a more viable option for extensive quantum optimization tasks, in contrast to QAOA, which may be more efficient for smaller-scale problems.
The remainder of the paper is organized as follows. In Section II, we introduce the formalism including definitions of QAOA and DC-QAOA. Section III provides the results. We show a comprehensive comparison of DC-QAOA and QAOA in Section III.1. In Section III.3, we discuss how DC-QAOA benefits from cd-terms, and in Section III.4, we demonstrate the parameter concentration phenomenon in DC-QAOA and propose a quantum resource-efficient training method. We conclude in Section IV.
@@ II Formalism
where $\lambda(t)\in[0,1]$ is the annealing schedule in $t\in[0,1]$. As $\dot{\lambda}\rightarrow 0$, the ground state of $H_{\mathrm{i}}$ can evolve adiabatically into the ground state of $H_{\mathrm{f}}$. Here, $H_{\mathrm{i}}=\sum_{k}h_{k}\sigma_{k}^{x}$, allowing the ground state of $H_{\mathrm{i}}$ to be easily prepared as an equally weighted superposition state in the computational bases, and $H_{\mathrm{f}}$ can be any Hamiltonian of interest, usually called problem Hamiltonian. Quantum annealing can be digitized with trotterized time evolution:
@@
However, the requirement for a large number of trotter layers $p$ in QAOA (sometimes referred to as “standard QAOA” hereafter) to approximate adiabatic evolution poses challenges due to the increased circuit complexity. To address this limitation, DC-QAOA has been developed Chandarana et al. (2022). DC-QAOA builds upon analog cd-driving which can shorten the evolution time of adiabatic transitions by adding an extra cd-term, presenting an opportunity to effectively reduce number of layers $p$ in QAOA. The nature of the cd-terms depends on the choice of appropriate adiabatic gauge potential Sels and Polkovnikov (2017). In the realm of analog cd-driving, the commonly utilized gauge potential is first-order nested commutators
@@
A schematic circuit diagram of a typical $p$-layer DC-QAOA is shown in Fig. 1. The detailed implementation of three possible cd-terms in DC-QAOA using elementary gates, including first-order nested commutator (hereafter referred to as NC) Chandarana et al. (2023), single qubit Y-rotation (referred to as Y) Čepaitė et al. (2023), and two-qubit YY-Interaction (referred to as YY) Chai et al. (2022), can be seen in Fig. 2.
@@ III.1 Circuit Length Scaling Behavior
For a MaxCut problem on graph $\mathcal{G}$ having $M$ edges, the implementation of $U_{\mathrm{f}}$ requires $2M$ CNOT gates, and the implementation of $U_{\mathrm{cd}}$ requires $4M$, $2M$ and 00 CNOT gates for the nested commutator (NC, Fig. 2(a)), YY-interaction (YY, Fig. 2(b)) and the single qubit Y-rotation (Y, Fig. 2(c)), respectively. Hereafter, we denote the three versions of the DC-QAOA algorithm which uses NC, YY and Y as cd-terms as DC-QAOA(NC), DC-QAOA(YY), and DC-QAOA(Y), respectively. Note that QAOA has $2M$ CNOT gates, therefore DC-QAOA(NC), DC-QAOA(YY), and DC-QAOA(Y) has $6M$, $4M$, and $2M$ CNOT gates in total respectively.
It is worth noting that DC-QAOA(NC) utilizes three times as many CNOT gates as QAOA, the most among the three versions of the algorithm. Such excessive usage of two-qubit gates in one layer is even more pronounced when we consider large-scale combinatorial optimization. For example, one-layer DC-QAOA(NC) requires $531867$ two-qubit gates while one-layer QAOA requires only $177289$ two-qubit gates based on the topology of Brisbane of IBM Quantum for a complete graph with 127 vertices, which is the number of vertices demonstrated in the most recent experiment Sachdeva et al. (2024). Therefore only assessing the complexity using number of layers is insufficient.
Device QAOA DC-QAOA(NC)
Brisbane Lalita et al. (2024) 177289 531867
Kyiv Tsunaki et al. (2024) 177634 532902
Torino Baglio (2024) 177944 533832
Table 1: The number of two-qubit gate of two algorithm when compiled into the topology of real quantum devices provided by IBM Quantum using Qiskit transpile function. The underlying graph of MaxCut problem is a complete graph. The native two-qubit gate being compiled to is Echoed Cross-Resonance Gate for Brisbane and Kyiv and Controlled-Z gate for Torino.
The performance of QAOA has been studied extensively in the context of combinatorial optimization problems, with scaling laws providing valuable insights into its efficiency. In Akshay et al. (2022), the authors proposed the logistic saturation conjecture in the context of MAX-2-SAT, demonstrating that the critical depth $p_{\mathrm{crit}}$, the minimum QAOA depth required to achieve a specific level of performance, follows a logistic function with respect to the clause density $\alpha$, This logistic model captures how $p_{\mathrm{crit}}$ initially grows rapidly with increasing $\alpha$, before saturating at a depth $p_{\mathrm{max}}$, beyond which further increases in circuit depth yield diminishing returns. Additionally, their study showed that $p_{\mathrm{max}}$ scales linearly with the problem size, suggesting that larger instances require proportionally deeper circuits.
In this work, we adapt these ideas to the MaxCut problem, where the analogous metric to clause density is the edge density of the graph. For a graph with $n$ vertices and $m$ edges, the edge density is defined as $\alpha_{\mathrm{e}}=\frac{2m}{n(n-1)}$, representing the fraction of possible edges that are present. We hypothesize that the logistic saturation conjecture also applies to MaxCut, with $p_{\mathrm{crit}}$ changing with the edge density of the graph following logistic function and then saturating at a depth $p_{\mathrm{max}}$. To support this, we provide a formal proof in Appendix A establishing the equivalence between Max-k-SAT and MaxCut. Moreover, we explore whether $p_{\mathrm{max}}$ retains a linear relationship with the graph size, providing further evidence for the generality of this scaling behavior across optimization problems. $p_{\mathrm{max}}$ provides a tool to quantify the advantage afforded by adding cd-terms in DC-QAOA. It is important to see how fast $p_{\mathrm{max}}$ scales with problem size. For example, if $p_{\mathrm{max}}$ of DC-QAOA(NC) scales three times slower than $p_{\mathrm{max}}$ of QAOA, it means that as the problem size grows, DC-QAOA(NC) can eventually solve an optimization problem utilizing fewer CNOT gates than QAOA. Therefore, it is fair to compare DC-QAOA and QAOA by comparing the scaling rate of $p_{\mathrm{max}}$ versus the problem size. To find the critical depth $p_{\mathrm{max}}$ for DC-QAOA and QAOA, we first define the critical depth in this study. In general, for a given graph with edge density $\alpha_{\mathrm{e}}$, the output state of the variational circuit will not cover the whole Hilbert space, the overlap, or fidelity $F$ between the optimal output state of a $p$-layer QAOA and the actual ground state |$\ket{\psi_{\mathrm{g}}}$ is defined as:
$F(p,\alpha_{\mathrm{e}})=\max_{\vec{\alpha},\vec{\beta}}\left|\bra{\psi_{\mathrm{g}}}U(\vec{\alpha},\vec{\beta})\ket{0}^{\otimes n}\right|^{2},$ (9)
and for a $p$-layer DC-QAOA, the fidelity is similarly defined as:
$F(p,\alpha_{\mathrm{e}})=\max_{\vec{\alpha},\vec{\beta},\vec{\gamma}}\left|\bra{\psi_{\mathrm{g}}}U(\vec{\alpha},\vec{\beta},\vec{\gamma})\ket{0}^{\otimes n}\right|^{2}.$ (10)
@@
Fig. 3 shows the fidelity $F$, calculated at fixed number of layers $p=20$, for a MaxCut problem with 16 vertices (cf. Fig. 12(h)) versus number of training iterations for four different versions of the algorithms. Different algorithms converge to different values of $F$, with a higher value indicating a better performance. Each panel shows eight color lines, which are 8 runs from random initialization of parameters. The converged value of $F$ for different algorithms are averaged and are shown in Table 2. Overall, the results from DC-QAOA(NC) (Fig. 3(b)) is the best, followed by those from DC-QAOA(YY) (Fig. 3(c)). Results from DC-QAOA(Y) (Fig. 3(d)) and QAOA (Fig. 3(a)) are comparable and are both inferior to the other two methods. These results suggest having more CNOT gates on each layer implies better performance of the algorithm.
@@
Figure 3: The fidelity $F$, calculated at fixed number of layers $p=20$, for a MaxCut problem versus number of training iterations for four different versions of the algorithms using same number of layers $p=20$. (a) QAOA, (b) DC-QAOA(NC), (c) DC-QAOA(YY) and (d) DC-QAOA(Y). The target graph has 16 vertices. In each panel, the 8 color lines represents 8 random initialization of parameters.
@@
Table 2: The average fidelity $F$ (“Mean”) for four different algorithms, along with their respective standard deviations (“Std.”).
Given some fixed tolerance on performance, $\epsilon>0$, for both algorithm on MaxCut problem defined on graph with edge density $\alpha_{\mathrm{e}}$, the critical depth $p_{\mathrm{crit}}$, required to guarantee approximations that fall within an $\epsilon>0$ tolerance, is given by:
$p_{\mathrm{crit}}(\alpha_{\mathrm{e}})=\min\{p|1-F(p,\alpha_{\mathrm{e}})<\epsilon\}.$ (11)
We can see in Fig. 4 that $p_{\mathrm{crit}}(\rho)$ of both QAOA and DC-QAOA(NC) in MaxCut problem also show an $S$-shaped curve, suggesting that logistic saturation conjecture still holds in this problem.
To elucidate how the combined effect of the number of CNOT gates and number of layers on the effectiveness of the algorithm, we focus on comparing DC-QAOA(NC), which has the most number of CNOT gates per layer, to QAOA.
Following the procedure in Akshay et al. (2022), we fit the $p_{\mathrm{crit}}(\rho)$ using the following logistic function:
$p_{\mathrm{crit}}(\alpha_{\mathrm{e}})=\frac{p_{\mathrm{max}}}{1+e^{k(\alpha_{\mathrm{e}}-\alpha_{\mathrm{e}}^{\mathrm{c}})}},$ (12)
@@
and in Fig. 5(a) we compare the saturating depth $p_{\mathrm{max}}$ as functions of the number of qubits $N$, for both QAOA and DC-QAOA(NC). We find that the saturation depth $p_{\mathrm{max}}$ also scales linearly with the graph size $N$ as shown in the Fig. 5(a). The calculated saturated depths $p_{\mathrm{max}}$ for both algorithms scale approximately linearly with system size as $p^{\mathrm{QAOA}}_{\mathrm{max}}\simeq 3.61N-19.01$ (for QAOA) and $p^{\mathrm{DC-QAOA(NC)}}_{\mathrm{max}}\simeq 0.88N-0.51$ (for DC-QAOA(NC)). Therefore, in the limit of solving large scale MaxCut problem on graph $\mathcal{G}$ with $M$ edges, QAOA needs $3.61N$ layers and in turn $3.61N\times 2M=7.22NM$ CNOT gates and DC-QAOA(NC) needs $0.82N$ layers and in turn $0.88N\times 6M=5.28NM$ CNOT gates. This implies that DC-QAOA(NC) is more advantageous than QAOA in solving large scale optimization problem because the total number of CNOT gates required to achieve the same level of effectiveness is smaller for DC-QAOA(NC), when $N$ is sufficiently large. In other words, the circuit complexity for DC-QAOA(NC) is lower, taking into consideration both the number of effective layers and the number of CNOT gates per layer required.
Figure 4: Average critical depth $p_{\mathrm{crit}}$ against edge density $\alpha_{\mathrm{e}}$ for QAOA in (a) and DC-QAOA(NC) in (b). Data points represent the average $p_{\mathrm{crit}}$ calculated according to Eq. (11) for 50 non-isomorphic graphs for all levels of edge density, and if the number of non-isomorphic graphs is less than 50, we make it up by choosing different initialization of both algorithm. Shaded area represent the standard deviation of $p_{\mathrm{crit}}$. Different color represent different system sizes.
@@
We set $3p^{\mathrm{QAOA}}_{\mathrm{max}}=p^{\mathrm{DC-QAOA(NC)}}_{\mathrm{max}}$ and obtain $N\simeq 16.54$. This suggests that DC-QAOA(NC) is more advantageous than QAOA when the number of vertices exceeds 16 (i.e. $N\geq 17$). To show this more clearly, we execute DC-QAOA(NC) with $p$ layers, and a QAOA with intentionally set $3p$ layers (which is denoted as QAOA-3 in Fig. 5(b),(c) and (d) and should not be confused with regular QAOA with $p$ layers), to solve the MaxCut problem with different number of vertices. Fig. 5(b) shows the results on 12 vertices (cf. Fig. 12(d)), and one can see that QAOA-3 clearly performs better than DC-QAOA(NC). Fig. 5(c) shows the results on 16 vertices (cf. Fig. 12(h)), in which situation the two algorithms are comparable. Given that $N=16$ is still less than 16.94, it is reasonable that QAOA-3 outperforms DC-QAOA(NC) slightly at this stage, although DC-QAOA(NC) rapidly approaches similar performance with increasing $p$. When $p>10$, the difference in $R$ between DC-QAOA(NC) and QAOA becomes negligible. We have also conducted a calculation comparing QAOA-3 and DC-QAOA(NC) for a problem with 20 vertices, with results shown in Fig. 5 (d), although the calculation is more expensive. In this case, DC-QAOA(NC) is clearly superior than QAOA-3. Our results shown in Fig. 5 confirm that DC-QAOA(NC) is only more advantageous than QAOA on graphs with more than 16 vertices. To our knowledge, this fact has not been noted in the literature and is the key result of this paper.
@@
Figure 5: (a) The saturating depth $p_{\mathrm{max}}$ versus the number of qubits $N$ for DC-QAOA(NC) and QAOA. (b)-(d) The fidelity to ground state $F$ versus $p$ for DC-QAOA(NC) with increasing $p$ layers and for QAOA with $3p$ layers (QAOA-3) using the same number of CNOT gates to find the maximum cut in a graph with (b) 12, (c) 16 and (d) 20 vertices.
III.2 Feasibility Test in NISQ devices
In the preceding section, we performed a comparative analysis of the DC-QAOA framework, with particular focus on its DC-QAOA(NC) variant, against the standard QAOA in a noiseless computational setting. This analysis sought to determine whether DC-QAOA(NC), which incorporates the counterdiabatic phenomenon, could surpass the conventional QAOA algorithm in performance as the size of the quantum system scales. A key feature of DC-QAOA(NC) is its reduced reliance on two-qubit gates, such as CNOT gates, which are known to be both error-prone and resource-intensive. By minimizing the usage of these gates, DC-QAOA(NC) presents a promising avenue for achieving greater efficiency and scalability compared to the traditional QAOA framework.
Nevertheless, it is essential to account for the practical limitations imposed by current quantum hardware, particularly in the NISQ era. Quantum circuits with substantial depth are often impractical due to the rapid onset of decoherence and other noise-induced challenges. As the circuit depth increases, errors from gate operations and interactions with the environment accumulate, significantly compromising the fidelity of quantum computations. These constraints necessitate quantum algorithms that balance performance with hardware feasibility. In this context, the practical value of DC-QAOA(NC) lies in its potential to address these limitations by reducing circuit depth, thus improving resilience to noise and decoherence. Examining the trade-offs between algorithmic efficiency and noise tolerance in realistic, hardware-constrained scenarios is vital to fully assessing the advantages of DC-QAOA(NC) over standard QAOA in the NISQ regime.
A widely adopted approach for evaluating the effectiveness of QAOA and its various extensions, both numerically and experimentally, involves benchmarking these algorithms under shallow-circuit conditions, typically with one or a few layers. Such configurations not only reduce the computational and experimental burden but also enable demonstrations on systems extending beyond a small number of qubits. These benchmarks provide critical insights into algorithmic performance and scalability while remaining compatible with the limitations of current quantum devices. Here, we benchmarked one-layer DC-QAOA(NC) and three-layer QAOA, which utilize same number of CNOT gates, on the Sherrington-Kirkpatrick (SK) model Sherrington and Kirkpatrick (1975). We calculate the fidelity between optimal output state of ansatz and the actual ground state, with 5 to 34 qubits in perfect environment using statevector simulator and with 5 to 17 qubits in noisy environment using quantum trajectory simulation with noise.
The SK model describes classical spin system with all-to-all couplings between the $n$ spins and its Hamiltonian is defined as follows:
$H_{\mathrm{SK}}=\frac{1}{2}\sum_{i,j}J_{ij}\sigma_{z}^{i}\sigma_{z}^{j},$ (13)
where $J_{ij}$ is sampled from standard normal distribution. In the noisy environment, each qubit suffers from depolarizing noise when undergoing the CNOT operation and two-qubit depolarization channel can be described by:
$\rho^{\prime}=(1-p)\rho+\frac{pI}{4}.$ (14)
In numerical simulation, such channel is applied after every two qubit gate, and by using quantum trajectory simulation method, we simulated this noisy quantum circuit up to 17 qubits. Here we set $p=0.66\%$, which is consistent with the CNOT error rate in the state-of-the-art devices published by IBM Quantum. The result of simulation of 100 randomly generated SK-model for every qubit $N$ can be seen from Fig. 6. Both Fig. 6(a) and Fig. 6(b) show a clear exponential trend with slowly increasing variance. In the perfect environment, the two algorithms reach the same level of performance at around $N=16$ which confirms that one-layer DC-QAOA(NC) has a significantly better exponential scaling compared to three-layer QAOA. Whereas in the noisy environment, the two algorithms reach the same level of performance at around $N=11$. This clearly shows that DC-QAOA(NC) is more robust compared to QAOA. This robustness can be attributed to the fact that counterdiabatic driving terms reduce complexity of the circuit. In Fig. 6(c) we can see that DC-QAOA(NC) shows exponential advantage over three-layer standard QAOA, and the advantage is more pronounced in the noisy environment.
Figure 6: Fidelity $F$ as a function of the number of qubits $N$ in the SK model, defined in Eq. (13). The comparison is made between a one-layer DC-QAOA(NC) and a three-layer QAOA, both utilizing the same number of CNOT gates. (a) Shows results in an ideal environment without noise, while (b) incorporates depolarization error. One-layer DC-QAOA(NC) performs better than three-layer after $N$ exceeds around 16 in an ideal environment while same happens when $N$ exceeds 11 in the noisy environment. In (c), we shows the exponential improvement of one-layer DC-QAOA(NC) over three-layer QAOA. In ideal environment, the improvement scales as $10^{0.0586N}$ and in noisy environment, the improvement scales as $10^{0.1397N}$
III.3 Effective Dimension and Excitation Suppression